OpenStudy (anonymous):

What is the solution set of the following system of equations? 5x - 3y - 10 = 0 x + y - 7 = 0

OpenStudy (anonymous):

5x-3y-10=0 =>x=(10+3y)/5 and, x+y-7=0 =>x=7-y ...............................................(a) Comparing x values, (10+3y)/5=7-y =>10+3y=5(7-y)=35-5y =>3y+5y=35-10 =>8y=25 =>y=25/8 Putting the value of y in equation (a), x=7-25/8=(56-25)/8=31/8 So the result is, (x,y)=(31/8,25/8)